(a) Find the solution of the initial-value problem and determine its interval of existence (i.e domain of the resulting function y).
(i) dy /dt = 1+t2 /t , y(t-1)=0
(ii) (t + 1)dy /dt = 1−y, y(t = 0) = 3. [Hint: Let A = −(±e−C).]
(b) Evaluate the following integrals: (i) xln(x +1)dx (ii) sin2(ln x)cos2(lnx)/ x dx.
1
Expert's answer
2020-07-02T19:29:45-0400

a)(i) Given "\frac{dy}{dt }= \frac{1+t^2}{ t}"

"\implies {dy}= (\frac{1}{ t}+t)dt"

Now by integrating, we get

"y=log(t)+\frac{t^2}{2} + c"

By given condition "y(1)=0"

"\implies c=-\frac{1}{2}"

Hence, solution is "y=log(t)+\frac{t^2}{2} -\frac{1}{2}" .

log(t) exists for postive values of t, hence interval of existence of solution is "(0,\infin)" .


a)(ii) Given "(t + 1)\frac{dy}{dt} = 1−y, y(t = 0) = 3"

"\implies \frac{dy}{1-y} =\frac{dt}{t+1}"

By integrating on both sides, we get

"-log(1-y) = log(t+1) -log(c)\\
\implies (1-y)(t+1) =c"

Now, by y(0)=3, "(1-3)(0+1)=c"

"\implies c=-2" .

Hence solution is "(1-y)(t+1)=-2
\implies y=1+\frac{2}{t+1}"

Solution y is not defined at t=-1 and solution contain y(t=0)=3.

Hence interval of existence of solution "(-1,\infin)" .


b)(i) Let I = "\u222b\n\nx\nln\n(\nx\n+\n1\n)\n \nd\nx" = "=
\frac{x^2}{2}
log
(
x
+
1
)
−
∫
 
\frac{x^2}{2}\frac{1}{x+1}
d
x"

Now for "\int \frac{x^2}{x+1}dx" , let "u = x+1 \implies dx = du"

So, "\int \frac{x^2}{x+1}dx = \int \frac{(u-1)^2}{u}du" = "\int (u-2+\frac{1}{u})du = \frac{u^2}{2} -2u+log(u) = \frac{(x+1)^2}{2} - 2(x+1) +log(x+1)" .

Hence, I = "\frac{x^2}{2}log(x+1)+\frac{(x+1)^2}{4} - (x+1) + \frac{log(x+1)}{2}"


b(ii) Let I = "\int \frac{sin (2(ln x))cos (2(lnx)) }{ x} dx"

Suppose "ln(x) = u \implies \frac{dx}{x} = du"

So, "\int \frac{sin (2(ln x))cos (2(lnx)) }{ x} dx = \int sin (2u)cos (2u) du"

"= \frac{1}{2} \int 2sin (2u)cos (2u) du = \frac{1}{2} \int sin(4u) du"

"= -\frac{1}{8} cos(4u) = -\frac{1}{8} cos(4 \ ln(x))"